how many 1's are between the numbers 1-500 ? I have tried to rite down all the numbers that have the number 1 in them and i came up with 150. Am I rite ? I am in 5th grade and this is for a study guide for a math tournament .
Answer by Fombitz(32388)

Let's do it this way.
From 000-009, there is 1 one(001).
From 010-019, there are 10+1=11 ones (011 counts as 2).
From 020-029, there is 1 one(021).
From 030-039, there is 1 one(031).
From 040-049, there is 1 one(041).
From 050-059, there is 1 one(051).
From 060-069, there is 1 one(061).
From 070-079, there is 1 one(071).
From 080-089, there is 1 one(081).
From 090-099, there is 1 one(091).
So from 000-099, there are 9+11=20 ones.
.
.
.
The same holds true for 200-299 (20),300-399 (20), 400-499 (20).
That's 80 so far.
.
.
.
It also holds for 100-199 but there is also the additional 1 in the hundreds place (20+100=120).
Let's add it all up.
80+120=200.
I count 200 ones.